OpenStudy (anonymous):

how do you take the derivative of y=ln(x^7)

OpenStudy (anonymous):

Chain rule = \[7x^6/x^7\] = \[7/x\]

OpenStudy (anonymous):

that's what i did, but it was marked wrong...

OpenStudy (unklerhaukus):

\[y=ln(x^7)\]\[y=7lnx\] \[y'=7 {1\over x }\]\[y'={7\over x}\]

OpenStudy (anonymous):

It's right. You're teacher is wrong. It happens :)

OpenStudy (anonymous):

Check this for "proof": http://www.wolframalpha.com/input/?i=derivative+ln%28x%5E7%29

OpenStudy (mimi_x3):

Use: \[\frac{d}{dx} \ln(...) = \frac{1}{(...)}*(...)' \] \[\frac{d}{dx} = \frac{1}{x^{7}} * 7x^6 = \frac{7x^{6}}{x^{7}} \]

OpenStudy (anonymous):

that was marked incorrectly as well... :-(

OpenStudy (mimi_x3):

Well, what about: \[\frac{7}{x} \]?

OpenStudy (anonymous):

i think my teacher is wrong...

OpenStudy (anonymous):

The derivative of \[\ln (x^7)\] has always been and always will be \[7/x\] :D

OpenStudy (anonymous):

it's 7/x

OpenStudy (anonymous):

You're teacher is DEFINITELY wrong. (Unless you wrote the question incorrectly)